# Provable Sample-Efficient Transfer Learning Conditional Diffusion Models via Representation Learning

### v1

@@ A.3 Proofs of Meta-Learning
$\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}\sim\mathbb{P}}\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{\widehat{f}^{\mathbb{P}},\widehat{h}})]$ (A.84)
@@
$\log\mathcal{N}_{\mathcal{F}}:=M_{f}^{2}S_{f}L_{f}\log\left(mnL_{f}W_{f}(B_{f}\vee 1)M_{f}T\log(1/\delta)\right),$ (A.85)
$\log\mathcal{N}_{\mathcal{H}}:=S_{h}L_{h}\log\left(nKL_{h}W_{h}(B_{h}\vee 1)M_{f}\gamma_{f}\log(1/\delta)\right).$ (A.86)
@@
$\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}\sim\mathbb{P}}\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{\widehat{f}^{\mathbb{P}},\widehat{h}})]$ (A.87)
@@
Here we apply Proposition A.4 in the first inequality, Proposition A.7 in the second and last inequality, Proposition A.5 in the fourth. Plugging in the definition of $r_{z},r_{P},r_{x}$ and $\log\mathcal{N}_{\mathcal{F}},\log\mathcal{N}_{\mathcal{H}}$ and noticing that $R_{f}\geq C_{R}\log^{\frac{1}{2}}(nKd_{x}M_{f}/\delta)\geq C_{R}^{\prime}\log^{\frac{1}{2}}\left(\frac{M_{f}K}{\log\mathcal{N}_{\mathcal{H}}}\right)$, we have with probability no less than $1-\delta$,
$\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathbb{E}_{\{(x_{i},y_{i})\}_{i=1}^{m}\sim\mathbb{P}}\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{\widehat{f}^{\mathbb{P}},\widehat{h}})]$ (A.88)
@@
A.4 Auxiliary Lemmas
Lemma A.9.
@@ A.4 Auxiliary Lemmas
$\mathcal{L}(\mathbb{P},h)\leq\mathbb{E}_{t,x_{t},y}\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}\leq C_{L}.$ (A.89)
@@
$\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{f,h})]$ $=\mathbb{E}_{(x,y)\sim\mathbb{P}}\mathbb{E}_{t,x_{t}|x}[\|f(x_{t},h(y),t)-\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}]$ (A.90)
@@
and $0\in\mathcal{F}$, it suffices to show that $\mathbb{E}_{t,x_{t},y}[\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}]$ is uniformly bounded for any $\mathbb{P},h$. According to (A.2),
$\mathbb{E}_{x_{t},y}[\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}]$ $\leq\mathbb{E}_{x_{t},y}\mathbb{E}_{x_{0}|(x_{t},y)}[\|\nabla_{x}\log\phi_{t}(x_{t}|x_{0})\|^{2}]$ (A.91)
@@
On the other hand, by (A.3) and Assumption 3.3,
$\mathbb{E}_{x_{t},y}[\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}]$ $\leq\mathbb{E}_{x_{t},y}\mathbb{E}_{x_{0}|(x_{t},y)}[\|\nabla_{x}\log p(x_{0}|y)\|^{2}\cdot e^{2t}]$ (A.92)
@@
$\mathcal{L}(\mathbb{P},h)$ $\leq\mathbb{E}_{t,x_{t},y}[\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}]$ (A.93)
@@
$\mathbb{E}_{x_{t}|x}\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}\leq C_{X}^{{}^{\prime\prime}}(\|x\|^{6}+1).$ (A.94)
@@
Note that $x_{t}|x\sim\mathcal{N}(x_{t}|\alpha_{t}x,\sigma_{t}^{2}I)$ and by Lemma A.2,
$\mathbb{E}_{x_{t}|x}\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}\leq\mathbb{E}_{x_{t}|x}2\Big{[}\|\nabla_{x}\log p_{t}(0|y)\|^{2}+(C_{X}+C_{X}^{\prime}\|x_{t}\|^{2})^{2}\|x_{t}\|^{2}\Big{]}$ (A.95)
@@
$\mathbb{E}_{q_{t}(x_{0}|0,y)}\|x_{0}\|^{2}\leq\mathbb{E}_{p(x_{0}|y)}\|x_{0}\|^{2}\leq C_{0}.$ (A.96)
According to (A.2),
$\|\nabla_{x}\log p_{t}(0|y)\|^{2}\leq\frac{\alpha_{t}^{2}}{\sigma_{t}^{4}}\mathbb{E}_{q_{t}(x_{0}|0,y)}\|x_{0}\|^{2}\leq\frac{C_{0}\alpha_{t}^{2}}{\sigma_{t}^{4}}.$ (A.97)
By (A.3), we also have
$\|\nabla_{x}\log p_{t}(0|y)\|^{2}\leq\frac{1}{\alpha_{t}^{2}}\mathbb{E}_{q_{t}(x_{0}|0,y)}\|\nabla_{x}\log p(x_{0}|y)\|^{2}\leq\frac{1}{\alpha_{t}^{2}}\mathbb{E}_{q_{t}(x_{0}|0,y)}[(B+L\|x_{0}\|)^{2}]\leq\frac{2(B^{2}+L^{2}C_{0})}{\alpha_{t}^{2}}.$ (A.98)
@@
$\|\nabla_{x}\log p_{t}(0|y)\|^{2}\leq(B^{2}+(L^{2}+1)C_{0})\cdot(\frac{\alpha_{t}^{2}}{\sigma_{t}^{4}}\wedge\frac{1}{\alpha_{t}^{2}})\leq 2(B^{2}+(L^{2}+1)C_{0}).$ (A.99)
Plug in (A.95) and we obtain for some constant $C_{X}^{{}^{\prime\prime}}$,
$\mathbb{E}_{x_{t}|x}\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}$ $\leq\mathbb{E}_{x_{t}|x}2\Big{[}(C_{X}+C_{X}^{\prime}\|x_{t}\|^{2})^{2}\|x_{t}\|^{2}\Big{]}+2(B^{2}+(L^{2}+1)C_{0})$ (A.100)
@@
Lemma A.11.
[…]
$\mathcal{R}_{n}(\Phi_{r}):=\mathbb{E}_{\bm{\sigma}}\sup_{\varphi\in\Phi_{r}}\Big{|}\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\varphi(x_{i})\Big{|}\leq\phi_{n}(r).$ (A.101)
@@
$\displaystyle\mathbb{E}_{\mathbb{P}}[\varphi]\leq\frac{2}{n}\sum_{i=1}^{n}\varphi(x_{i})+C^{\prime}(B\vee b)\left(r_{n}^{*}+\frac{\log\big{(}(\log n)/\delta\big{)}}{n}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big{(}(\log n)/\delta\big{)}}{n}},$ (A.102)
$\displaystyle\frac{1}{n}\sum_{i=1}^{n}\varphi(x_{i})\leq 2\mathbb{E}_{\mathbb{P}}[\varphi]+C^{\prime}(B\vee b)\left(r_{n}^{*}+\frac{\log\big{(}(\log n)/\delta\big{)}}{n}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big{(}(\log n)/\delta\big{)}}{n}}.$ (A.103)
@@
$\Phi^{(j)}:=\{\varphi\in\Phi:\epsilon_{j+1}<\mathbb{E}_{\mathbb{P}}[\varphi]\leq\epsilon_{j}\}.$ (A.104)
@@
$\displaystyle\Big{|}\frac{1}{n}\sum_{i=1}^{n}\varphi(x_{i})-\mathbb{E}_{\mathbb{P}}[\varphi]\Big{|}\lesssim\mathcal{R}_{n}(\Phi^{(j)})+\sqrt{\frac{(B\epsilon_{j}+B_{0})\log\big{(}\log(b/\epsilon_{j})/\delta\big{)}}{n}}+\frac{b\log\big{(}\log(b/\epsilon_{j})/\delta\big{)}}{n},$ (A.105)
$\displaystyle\Big{|}\frac{1}{n}\sum_{i=1}^{n}(\varphi(x_{i}))^{2}-\mathbb{E}_{\mathbb{P}}[\varphi^{2}]\Big{|}\lesssim b\mathcal{R}_{n}(\Phi^{(j)})+\sqrt{\frac{b^{2}(B\epsilon_{j}+B_{0})\log\big{(}\log(b/\epsilon_{j})/\delta\big{)}}{n}}+\frac{b^{2}\log\big{(}\log(b/\epsilon_{j})/\delta\big{)}}{n}.$ (A.106)
@@
$\Big{|}\frac{1}{n}\sum_{i=1}^{n}\varphi(x_{i})-\mathbb{E}_{\mathbb{P}}[\varphi]\Big{|}\lesssim\mathcal{R}_{n}(\Phi^{(j_{0}:)})+\sqrt{\frac{(B\epsilon_{j_{0}}+B_{0})\log\big{(}\log(n)/\delta\big{)}}{n}}+\frac{b\log\big{(}(\log n)/\delta\big{)}}{n}$ (A.107)
From now on we reason on the conjunction of (A.105), (A.106) and (A.107). Define
$U_{j}=B\epsilon_{j}+B_{0}+b\mathcal{R}_{n}(\Phi^{(k)})+\sqrt{\frac{b^{2}(B\epsilon_{j}+B_{0})\log\big{(}\log(b/\epsilon_{j})/\delta\big{)}}{n}}+\frac{b^{2}\log\big{(}\log(b/\epsilon_{j})/\delta\big{)}}{n}.$ (A.108)
and thus for any $\varphi\in\Phi^{(j)}$, we have $\frac{1}{n}\sum_{i=1}^{n}(\varphi(x_{i}))^{2}\leq CU_{j}$ for some absolute constant $C$ by (A.106), indicating that $\mathcal{R}_{n}(\Phi^{(j)})\leq\phi_{n}(CU_{j})\leq\sqrt{C}\phi_{n}(U_{j})$. For any $j\leq j_{0}$,
$U_{j}\leq 2(B\epsilon_{j}+B_{0})+b\sqrt{C}\phi_{n}(U_{j})+\frac{2b^{2}\log\big{(}(\log n)/\delta\big{)}}{n}.$ (A.109)
@@
$U_{j}\lesssim b^{2}r_{n}^{*}+B\epsilon_{j}+B_{0}+\frac{b^{2}\log\big{(}(\log n)/\delta\big{)}}{n}=:r_{n}(\epsilon_{j}).$ (A.110)
Therefore, for any $\varphi\in\Phi^{(j)},j\leq j_{0}$, by (A.105),
$\Big{|}\frac{1}{n}\sum_{i=1}^{n}\varphi(x_{i})-\mathbb{E}_{\mathbb{P}}[\varphi]\Big{|}$ $\lesssim\phi_{n}(r_{n}(\epsilon_{j}))+\sqrt{\frac{(B\epsilon_{j}+B_{0})\log\big{(}(\log n)/\delta\big{)}}{n}}+\frac{b\log\big{(}(\log n)/\delta\big{)}}{n}$ (A.111)
@@
$\Big{|}\frac{1}{n}\sum_{i=1}^{n}\varphi(x_{i})-\mathbb{E}_{\mathbb{P}}[\varphi]\Big{|}\lesssim F_{n}(\mathbb{E}_{\mathbb{P}}[\varphi]).$ (A.112)
@@
$\displaystyle\mathbb{E}_{\mathbb{P}}[\varphi]\leq\frac{2}{n}\sum_{i=1}^{n}\varphi(x_{i})+C^{\prime}(B\vee b)\left(r_{n}^{*}+\frac{\log\big{(}(\log n)/\delta\big{)}}{n}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big{(}(\log n)/\delta\big{)}}{n}},$ (A.113)
$\displaystyle\frac{1}{n}\sum_{i=1}^{n}\varphi(x_{i})\leq 2\mathbb{E}_{\mathbb{P}}[\varphi]+C^{\prime}(B\vee b)\left(r_{n}^{*}+\frac{\log\big{(}(\log n)/\delta\big{)}}{n}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big{(}(\log n)/\delta\big{)}}{n}}.$ (A.114)
Here $C^{\prime}$ is an absolute constant. Moreover, when $\varphi\in\Phi^{(j)}$ for $j>j_{0}$, we have $\mathbb{E}_{\mathbb{P}}[\varphi]\leq\frac{b}{n}$, and according to (A.107),
$\Big{|}\frac{1}{n}\sum_{i=1}^{n}\varphi(x_{i})-\mathbb{E}_{\mathbb{P}}[\varphi]\Big{|}\lesssim F_{n}(\varepsilon_{j_{0}}).$ (A.115)
@@
Lemma A.12.
[…]
$\mathcal{R}_{K,n}(\Phi_{r}):=\mathbb{E}_{\bm{\sigma}}\sup_{\varphi\in\Phi_{r}}\Big{|}\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\sigma_{i}^{k}\varphi(x_{i}^{k})\Big{|}\leq\phi_{K,n}(r).$ (A.116)
@@
$\displaystyle\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\leq\frac{2}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i})+C^{\prime}(B\vee b)\left(r_{K,n}^{*}+\frac{\log\big{(}(\log nK)/\delta\big{)}}{nK}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big{(}(\log nK)/\delta\big{)}}{nK}},$ (A.117)
$\displaystyle\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})\leq 2\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]+C^{\prime}(B\vee b)\left(r_{K,n}^{*}+\frac{\log\big{(}(\log nK)/\delta\big{)}}{nK}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big{(}(\log nK)/\delta\big{)}}{nK}}.$ (A.118)
@@
$\Phi^{(j)}:=\{\varphi\in\Phi:\epsilon_{j+1}<\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\leq\epsilon_{j}\}.$ (A.119)

### v2

@@ B.3 Proofs of Meta-Learning
[…]
@@
$\log\mathcal{N}_{\mathcal{F}}:=M_{f}^{2}S_{f}L_{f}\log\left(mnL_{f}W_{f}(B_{f}\vee 1)M_{f}T\log(1/\delta)\right),$ (B.85)
$\log\mathcal{N}_{\mathcal{H}}:=S_{h}L_{h}\log\left(nKL_{h}W_{h}(B_{h}\vee 1)M_{f}\gamma_{f}\log(1/\delta)\right).$ (B.86)
@@
[…]
@@
Here we apply Proposition B.4 in the first inequality, Proposition B.7 in the second and last inequality, Proposition B.5 in the fourth. Plugging in the definition of $r_{z},r_{P},r_{x}$ and $\log\mathcal{N}_{\mathcal{F}},\log\mathcal{N}_{\mathcal{H}}$ and noticing that $R_{f}\geq C_{R}\log^{\frac{1}{2}}(nKd_{x}M_{f}/\delta)\geq C_{R}^{\prime}\log^{\frac{1}{2}}\left(\frac{M_{f}K}{\log\mathcal{N}_{\mathcal{H}}}\right)$, we have with probability no less than $1-\delta$,
[…]
@@
B.4 Auxiliary Lemmas
Lemma B.9.
@@ B.4 Auxiliary Lemmas
$\mathcal{L}(\mathbb{P},h)\leq\mathbb{E}_{t,x_{t},y}\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}\leq C_{L}.$ (B.89)
@@
$\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{f,h})]$ $=\mathbb{E}_{(x,y)\sim\mathbb{P}}\mathbb{E}_{t,x_{t}|x}[\|f(x_{t},h(y),t)-\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}]$ (B.90)
@@
and $0\in\mathcal{F}$, it suffices to show that $\mathbb{E}_{t,x_{t},y}[\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}]$ is uniformly bounded for any $\mathbb{P},h$. According to (B.2),
$\mathbb{E}_{x_{t},y}[\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}]$ $\leq\mathbb{E}_{x_{t},y}\mathbb{E}_{x_{0}|(x_{t},y)}[\|\nabla_{x}\log\phi_{t}(x_{t}|x_{0})\|^{2}]$ (B.91)
@@
On the other hand, by (B.3) and Assumption 3.3,
$\mathbb{E}_{x_{t},y}[\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}]$ $\leq\mathbb{E}_{x_{t},y}\mathbb{E}_{x_{0}|(x_{t},y)}[\|\nabla_{x}\log p(x_{0}|y)\|^{2}\cdot e^{2t}]$ (B.92)
@@
$\mathcal{L}(\mathbb{P},h)$ $\leq\mathbb{E}_{t,x_{t},y}[\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}]$ (B.93)
@@
$\mathbb{E}_{x_{t}|x}\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}\leq C_{X}^{{}^{\prime\prime}}(\|x\|^{6}+1).$ (B.94)
@@
Note that $x_{t}|x\sim\mathcal{N}(x_{t}|\alpha_{t}x,\sigma_{t}^{2}I)$ and by Lemma B.2,
$\mathbb{E}_{x_{t}|x}\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}\leq\mathbb{E}_{x_{t}|x}2\Big[\|\nabla_{x}\log p_{t}(0|y)\|^{2}+(C_{X}+C_{X}^{\prime}\|x_{t}\|^{2})^{2}\|x_{t}\|^{2}\Big]$ (B.95)
@@
$\mathbb{E}_{q_{t}(x_{0}|0,y)}\|x_{0}\|^{2}\leq\mathbb{E}_{p(x_{0}|y)}\|x_{0}\|^{2}\leq C_{0}.$ (B.96)
According to (B.2),
$\|\nabla_{x}\log p_{t}(0|y)\|^{2}\leq\frac{\alpha_{t}^{2}}{\sigma_{t}^{4}}\mathbb{E}_{q_{t}(x_{0}|0,y)}\|x_{0}\|^{2}\leq\frac{C_{0}\alpha_{t}^{2}}{\sigma_{t}^{4}}.$ (B.97)
By (B.3), we also have
$\|\nabla_{x}\log p_{t}(0|y)\|^{2}\leq\frac{1}{\alpha_{t}^{2}}\mathbb{E}_{q_{t}(x_{0}|0,y)}\|\nabla_{x}\log p(x_{0}|y)\|^{2}\leq\frac{1}{\alpha_{t}^{2}}\mathbb{E}_{q_{t}(x_{0}|0,y)}[(B+L\|x_{0}\|)^{2}]\leq\frac{2(B^{2}+L^{2}C_{0})}{\alpha_{t}^{2}}.$ (B.98)
@@
$\|\nabla_{x}\log p_{t}(0|y)\|^{2}\leq(B^{2}+(L^{2}+1)C_{0})\cdot(\frac{\alpha_{t}^{2}}{\sigma_{t}^{4}}\wedge\frac{1}{\alpha_{t}^{2}})\leq 2(B^{2}+(L^{2}+1)C_{0}).$ (B.99)
Plug in (B.95) and we obtain for some constant $C_{X}^{{}^{\prime\prime}}$,
$\mathbb{E}_{x_{t}|x}\|\nabla_{x}\log p_{t}(x_{t}|y)\|^{2}$ $\leq\mathbb{E}_{x_{t}|x}2\Big[(C_{X}+C_{X}^{\prime}\|x_{t}\|^{2})^{2}\|x_{t}\|^{2}\Big]+2(B^{2}+(L^{2}+1)C_{0})$ (B.100)
@@
Lemma B.11.
Let $\Phi$ be a class of functions on domain $\Omega$ and $\mathbb{P}$ be a probability distribution over $\Omega$. Suppose that for any $\varphi\in\Phi$, $\|\varphi\|_{L^{\infty}(\Omega)}\leq b$, $\mathbb{E}_{\mathbb{P}}[\varphi]\geq 0$, and $\mathbb{E}_{\mathbb{P}}[\varphi^{2}]\leq B\mathbb{E}_{\mathbb{P}}[\varphi]+B_{0}$ for some $b,B,B_{0}\geq 0$. Let $x_{1},\cdots,x_{n}\overset{\textit{i.i.d.}}{\sim}\mathbb{P}$ and $\phi_{n}$ be a positive, non-decreasing and sub-root function such that
$\mathcal{R}_{n}(\Phi_{r}):=\mathbb{E}_{\bm{\sigma}}\sup_{\varphi\in\Phi_{r}}\Big|\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\varphi(x_{i})\Big|\leq\phi_{n}(r).$ (B.101)
@@
$\displaystyle\mathbb{E}_{\mathbb{P}}[\varphi]\leq\frac{2}{n}\sum_{i=1}^{n}\varphi(x_{i})+C^{\prime}(B\vee b)\left(r_{n}^{*}+\frac{\log\big((\log n)/\delta\big)}{n}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big((\log n)/\delta\big)}{n}},$ (B.102)
$\displaystyle\frac{1}{n}\sum_{i=1}^{n}\varphi(x_{i})\leq 2\mathbb{E}_{\mathbb{P}}[\varphi]+C^{\prime}(B\vee b)\left(r_{n}^{*}+\frac{\log\big((\log n)/\delta\big)}{n}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big((\log n)/\delta\big)}{n}}.$ (B.103)
@@
$\Phi^{(j)}:=\{\varphi\in\Phi:\epsilon_{j+1}<\mathbb{E}_{\mathbb{P}}[\varphi]\leq\epsilon_{j}\}.$ (B.104)
@@
$\displaystyle\Big|\frac{1}{n}\sum_{i=1}^{n}\varphi(x_{i})-\mathbb{E}_{\mathbb{P}}[\varphi]\Big|\lesssim\mathcal{R}_{n}(\Phi^{(j)})+\sqrt{\frac{(B\epsilon_{j}+B_{0})\log\big(\log(b/\epsilon_{j})/\delta\big)}{n}}+\frac{b\log\big(\log(b/\epsilon_{j})/\delta\big)}{n},$ (B.105)
$\displaystyle\Big|\frac{1}{n}\sum_{i=1}^{n}(\varphi(x_{i}))^{2}-\mathbb{E}_{\mathbb{P}}[\varphi^{2}]\Big|\lesssim b\mathcal{R}_{n}(\Phi^{(j)})+\sqrt{\frac{b^{2}(B\epsilon_{j}+B_{0})\log\big(\log(b/\epsilon_{j})/\delta\big)}{n}}+\frac{b^{2}\log\big(\log(b/\epsilon_{j})/\delta\big)}{n}.$ (B.106)
@@
$\Big|\frac{1}{n}\sum_{i=1}^{n}\varphi(x_{i})-\mathbb{E}_{\mathbb{P}}[\varphi]\Big|\lesssim\mathcal{R}_{n}(\Phi^{(j_{0}:)})+\sqrt{\frac{(B\epsilon_{j_{0}}+B_{0})\log\big(\log(n)/\delta\big)}{n}}+\frac{b\log\big((\log n)/\delta\big)}{n}$ (B.107)
From now on we reason on the conjunction of (B.105), (B.106) and (B.107). Define
$U_{j}=B\epsilon_{j}+B_{0}+b\mathcal{R}_{n}(\Phi^{(k)})+\sqrt{\frac{b^{2}(B\epsilon_{j}+B_{0})\log\big(\log(b/\epsilon_{j})/\delta\big)}{n}}+\frac{b^{2}\log\big(\log(b/\epsilon_{j})/\delta\big)}{n}.$ (B.108)
and thus for any $\varphi\in\Phi^{(j)}$, we have $\frac{1}{n}\sum_{i=1}^{n}(\varphi(x_{i}))^{2}\leq CU_{j}$ for some absolute constant $C$ by (B.106), indicating that $\mathcal{R}_{n}(\Phi^{(j)})\leq\phi_{n}(CU_{j})\leq\sqrt{C}\phi_{n}(U_{j})$. For any $j\leq j_{0}$,
$U_{j}\leq 2(B\epsilon_{j}+B_{0})+b\sqrt{C}\phi_{n}(U_{j})+\frac{2b^{2}\log\big((\log n)/\delta\big)}{n}.$ (B.109)
@@
$U_{j}\lesssim b^{2}r_{n}^{*}+B\epsilon_{j}+B_{0}+\frac{b^{2}\log\big((\log n)/\delta\big)}{n}=:r_{n}(\epsilon_{j}).$ (B.110)
Therefore, for any $\varphi\in\Phi^{(j)},j\leq j_{0}$, by (B.105),
$\Big|\frac{1}{n}\sum_{i=1}^{n}\varphi(x_{i})-\mathbb{E}_{\mathbb{P}}[\varphi]\Big|$ $\lesssim\phi_{n}(r_{n}(\epsilon_{j}))+\sqrt{\frac{(B\epsilon_{j}+B_{0})\log\big((\log n)/\delta\big)}{n}}+\frac{b\log\big((\log n)/\delta\big)}{n}$ (B.111)
@@
$\Big|\frac{1}{n}\sum_{i=1}^{n}\varphi(x_{i})-\mathbb{E}_{\mathbb{P}}[\varphi]\Big|\lesssim F_{n}(\mathbb{E}_{\mathbb{P}}[\varphi]).$ (B.112)
@@
$\displaystyle\mathbb{E}_{\mathbb{P}}[\varphi]\leq\frac{2}{n}\sum_{i=1}^{n}\varphi(x_{i})+C^{\prime}(B\vee b)\left(r_{n}^{*}+\frac{\log\big((\log n)/\delta\big)}{n}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big((\log n)/\delta\big)}{n}},$ (B.113)
$\displaystyle\frac{1}{n}\sum_{i=1}^{n}\varphi(x_{i})\leq 2\mathbb{E}_{\mathbb{P}}[\varphi]+C^{\prime}(B\vee b)\left(r_{n}^{*}+\frac{\log\big((\log n)/\delta\big)}{n}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big((\log n)/\delta\big)}{n}}.$ (B.114)
Here $C^{\prime}$ is an absolute constant. Moreover, when $\varphi\in\Phi^{(j)}$ for $j>j_{0}$, we have $\mathbb{E}_{\mathbb{P}}[\varphi]\leq\frac{b}{n}$, and according to (B.107),
$\Big|\frac{1}{n}\sum_{i=1}^{n}\varphi(x_{i})-\mathbb{E}_{\mathbb{P}}[\varphi]\Big|\lesssim F_{n}(\varepsilon_{j_{0}}).$ (B.115)
@@
Lemma B.12.
Let $\Phi$ be a class of functions on domain $\Omega$, $\mathbb{P}^{1},\cdots,\mathbb{P}^{K}$ be probability distributions over $\Omega$, and $\widehat{\mathbb{P}}^{(K)}=\frac{1}{K}\sum_{k=1}^{K}\delta_{\mathbb{P}^{k}}$. Suppose that for any $\varphi\in\Phi$, $\|\varphi\|_{L^{\infty}(\Omega)}\leq b$, $\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\geq 0$, and $\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi^{2}]\leq B\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]+B_{0}$ for some $b,B,B_{0}\geq 0$. Let $x^{k}_{1},\cdots,x^{k}_{n}\overset{\textit{i.i.d.}}{\sim}\mathbb{P}^{k}$ for any $k$ and all $(x_{i}^{k})_{i,k}$ are independent. Let $\phi_{K,n}$ be a positive, non-decreasing and sub-root function such that
$\mathcal{R}_{K,n}(\Phi_{r}):=\mathbb{E}_{\bm{\sigma}}\sup_{\varphi\in\Phi_{r}}\Big|\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\sigma_{i}^{k}\varphi(x_{i}^{k})\Big|\leq\phi_{K,n}(r).$ (B.116)
@@
$\displaystyle\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\leq\frac{2}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i})+C^{\prime}(B\vee b)\left(r_{K,n}^{*}+\frac{\log\big((\log nK)/\delta\big)}{nK}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big((\log nK)/\delta\big)}{nK}},$ (B.117)
$\displaystyle\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})\leq 2\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]+C^{\prime}(B\vee b)\left(r_{K,n}^{*}+\frac{\log\big((\log nK)/\delta\big)}{nK}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big((\log nK)/\delta\big)}{nK}}.$ (B.118)
@@
$\Phi^{(j)}:=\{\varphi\in\Phi:\epsilon_{j+1}<\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\leq\epsilon_{j}\}.$ (B.119)